½(y-2)+2 =3/8(3y-4)
How exactly do you solve for a problem like this?? i get 7x=7x i dont think this answer is right

Answer by stanbon(75887)

½(y-2)+2 =3/8(3y-4)
---
Multiply both sides by 8 to get:
4(y-2) + 16 = 3(3y-4)
---
4y-8 + 16 = 9y-12
---
4y+8 = 9y-12
5y = 20
y = 4
------------
Cheers,
Stan H.
===============